OpenStudy (anonymous):

determine weather the point(1,8) is a solution of the system by substituting the point into each eqution, 2x+6y=50, 10x+0y=10

OpenStudy (anonymous):

so for the first equation: 2(1)+6(8)=50 2+48=50 50=50, so yes, it is a solution for the second: 10(1)+0(8)=10 10+0=10, so again, yes this is a solution